# Endmember Extraction from Hyperspectral Images Using Self-Dictionary Approach with Linear Programming

### v2

@@ VII Concluding Remarks
This study was motivated by the following question: although the theoretical results shown in [16, 28] suggest that Hottopixx methods are effective for endmember extraction problems, is this really true? Hottopixx methods require solving computationally expensive LP problems, called Hottopixx models, which makes them challenging to implement in practice. To address this issue, we developed an efficient and effective implementation of Hottopixx, EEHT, and found numerical evidence indicating that EEHT can estimate the endmember signatures of real HSIs with reasonable accuracy.
@@
We close this paper with remarks on potential directions for future research. As we saw in Section VI-C, EEHT can be applied to endmember extraction of HSIs with around one hundred thousand pixels. However, its computational time is much longer than that of greedy methods, such as SPA and its refinements. In order to reduce it, we need to focus on the computational cost of RCE, whose bottleneck lies in computing the optimal solutions to problems $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ in step 2-1, since the problem size can be large. Augmented Lagrangian methods are often used for solving large-scale LPs; see [8], for instance. These methods may be of help in reducing the computational time required by EEHT.
@@ Appendix A Proof of Theorem 2
First, we show that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))\geq\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{N}))$ holds under condition (C1). The matrix $[X^{*},\Gamma^{*}]$ belongs to $\mathcal{F}(\ell,n)$, and thus, it is a feasible solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$. In addition, we obtain $\|A(\mathcal{L})-A(\mathcal{L})X^{*}\|_{1}=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by Theorem 1 (ii), and $\|A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}]\|_{1}\leq\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by condition (C1) and Theorem 1 (i). Thus, the objective function value of $\mathsf{H}(\mathcal{L},\mathcal{N})$ at $[X^{*},\Gamma^{*}]$ satisfies
@@
Next, we show the reverse inequality. Let $\hat{X}^{*}=[\hat{\bm{x}}_{1}^{*},\ldots,\hat{\bm{x}}_{n}^{*}]$ denote the optimal solution of $\mathsf{H}(\mathcal{L},\mathcal{N})$. We find that
@@
The last inequality comes from that the matrix $[\hat{\bm{x}}_{1}^{*},\ldots,\hat{\bm{x}}_{\ell}^{*}]$ belongs to $\mathcal{F}(\ell,\ell)$ and is, therefore, a feasible solution to $\mathsf{H}(\mathcal{L},\mathcal{L})$. The two inequalities we showed above yield
@@
We conclude that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{N}))$ and the matrix $[X^{*},\Gamma^{*}]$ is the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$. ∎
@@
  • Lemma 2: Let $\alpha^{*}$ be the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{L})$. If $\alpha^{*}$ satisfies condition (C1), then $\alpha$ can be constructed from $\alpha^{*}$ such that $\alpha$ is a feasible solution to $\mathsf{P}(\mathcal{N},\mathcal{N})$, and the objective function value of $\mathsf{P}(\mathcal{N},\mathcal{N})$ at $\alpha$ is $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$.
  • Lemma 3: Let $\beta^{*}$ be the optimal solution of $\mathsf{D}(\mathcal{L},\mathcal{L})$. If $\beta^{*}$ satisfies condition (C2), then $\beta$ can be constructed from $\beta^{*}$ such that $\beta$ is a feasible solution of $\mathsf{D}(\mathcal{N},\mathcal{N})$, and the objective function value of $\mathsf{D}(\mathcal{N},\mathcal{N})$ at $\beta$ is $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))$.
Theorem 1 (i) tells us that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))$. Accordingly, as we reviewed in Section II-B, the weak duality theorem implies that $\alpha$ and $\beta$ are optimal solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$, respectively. Furthermore, it means that $\mathrm{opt}(\mathsf{P}(\mathcal{N},\mathcal{N}))=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. As a result, using parts (i) and (ii) of Theorem 1, we can establish part (ii) of Theorem 2.
@@
Let $\mathcal{L}\subset\mathcal{N}$. Let $\alpha^{*}=(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution of $\mathsf{P}(\mathcal{L},\mathcal{L})$. Assume that condition (C1) holds. By letting $\Gamma^{*}=[\bm{\gamma}_{j}^{*}:j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}_{j}^{*}$ of $\mathsf{R}_{j}(\mathcal{L},X^{*})$ and letting $\Pi$ be a permutation matrix of size $n$ satisfying $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$, construct $\alpha=(X,F,G,u)\in\mathbb{R}^{n\times n}\times\mathbb{R}^{d\times n}\times\mathbb{R}^{d\times n}\times\mathbb{R}$ as follows:
@@
where $R=A-AX$ for $X$ as constructed above. Then, the following hold.
@@
Part (i)  It is sufficient to prove that $\alpha$ satisfies the first, second, and fifth constraints of $\mathsf{P}(\mathcal{N},\mathcal{N})$, since $F=R^{+}\geq O$ and $G=R^{-}\geq O$. Let us look at the first constraint. Since $R$, $F$ and $G$ are constructed as $R=A-AX$, $F=R^{+}$ and $G=R^{-}$, we have
@@
for $S=A(\mathcal{L})-A(\mathcal{L})X^{*}$ and $T=A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}$. It follows from parts (i) and (ii) of Theorem 1 that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\|A(\mathcal{L})-A(\mathcal{L})X^{*}\|_{1}=\|S\|_{1}$. It follows from condition (C1) that $\|T\|_{1}=\|A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}\|_{1}\leq\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. Consequently, we obtain $u=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. ∎
@@
Let $\mathcal{L}\subset\mathcal{N}$. Let $\beta^{*}=(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution of $\mathsf{D}(\mathcal{L},\mathcal{L})$. Assume that condition (C2) holds. By letting $\Pi$ be a permutation matrix of size $n$ satisfying $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$, and $\Delta=(Y^{*})^{\top}A(\mathcal{N}\setminus\mathcal{L})$, construct $\beta=(Y,Z,\bm{s},\bm{t},v)\in\mathbb{R}^{d\times n}\times\mathbb{R}^{n\times n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$ as follows:
@@
$\Leftrightarrow$ $-J\cdot\mathrm{diag}(\Pi\hat{\bm{s}})\leq\hat{Y}\Pi^{\top}\leq J\cdot\mathrm{diag}(\Pi\hat{\bm{s}})$
@@
$\Leftrightarrow$ $-J\cdot\mathrm{diag}(\hat{\bm{s}})\leq\hat{Y}\leq J\cdot\mathrm{diag}(\hat{\bm{s}})$
@@
$\mathrm{diag}(\hat{\bm{t}})$ $=\begin{bmatrix}\mathrm{diag}(\bm{t}^{*})&O\\ O&O\end{bmatrix},$
@@
Let $\alpha$ and $\beta$ be as constructed in Lemmas 2 and 3. From the lemmas and Theorem 1 (i), we find that $\alpha$ and $\beta$ are feasible solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$ and the objective function value of $\mathsf{P}(\mathcal{N},\mathcal{N})$ at $\alpha$ is equal to that of $\mathsf{D}(\mathcal{N},\mathcal{N})$ at $\beta$. Thus, the weak duality theorem implies that $\alpha$ and $\beta$ are optimal solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$. Furthermore, this means that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{P}(\mathcal{N},\mathcal{N}))$. Consequently, it follows from parts (i) and (ii) of Theorem 1 that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{N},\mathcal{N}))$ holds and $X$ of $\alpha$ is the optimal solution to $\mathsf{H}(\mathcal{N},\mathcal{N})$. ∎
Appendix B Cluster Construction in Step 2-1 of Algorithm 1
Here, we describe the notation for method of constructing a cluster $\mathcal{S}_{\ell}$ in step 2-1 of Algorithm 1. We choose a column $\bm{a}_{i}$ of $A$ and sort the columns $\bm{a}_{1},\ldots,\bm{a}_{n}$ of $A$ by their distance to $\bm{a}_{i}$ in ascending order so that

### v3

@@ VIII Concluding Remarks
This study was motivated by the following question: although the theoretical results shown in [17, 31] suggest that Hottopixx methods are effective for endmember extraction problems, is this really true? Hottopixx methods require solving computationally expensive LP problems, called Hottopixx models, which makes them challenging to use in practice. To address this issue, we developed an efficient and effective implementation of Hottopixx, EEHT, and provided numerical evidence indicating that EEHT can estimate the endmember signatures of real HSIs with reasonable accuracy.
We conclude this paper with remarks on potential directions for future research. Further theoretical investigation is needed to clarify the computational efficiency of RCE, which is incorporated in EEHT. In particular, it is important to analyze its iteration complexity as well as the growth of the index set $\mathcal{L}$ during the iterations.
@@
As we saw in Section VII-C, EEHT can be applied to endmember extraction of HSIs with on the order of one hundred thousand pixels. However, its computational time is considerably longer than that of greedy methods, such as SPA and its refinements. To reduce this, we need to focus on the computational cost of RCE, whose bottleneck lies in solving problems $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$ in step 2-1, since the problem size can be large. Augmented Lagrangian methods are often employed for solving large-scale LPs; see [8], for instance. These methods may help reduce the computational time required by EEHT.
Another promising direction is the use of preprocessing techniques to remove redundant pixels from HSI data that are far from pure pixels. Such techniques can substantially reduce the computational burden of solving $\mathsf{P}(\mathcal{L},\mathcal{L})$ and $\mathsf{D}(\mathcal{L},\mathcal{L})$. For this approach to be successful, it is important to develop preprocessing methods that ensure the remaining pixels still contain pure pixels or those close to them. We are currently pursuing research along this line and expect to report our findings in the near future.
@@ Appendix A Proof of Theorem 1
Proof:
Part (i)  We show that feasible solutions exist for both $\mathsf{P}(\mathcal{L},\mathcal{M})$ and $\mathsf{D}(\mathcal{L},\mathcal{M})$. Setting all variables of $\mathsf{D}(\mathcal{L},\mathcal{M})$ to zero yields a solution that satisfies all of its constraints. Thus, $\mathsf{D}(\mathcal{L},\mathcal{M})$ has a feasible solution. Moreover, $\mathcal{F}(\ell,m)$ is nonempty. Indeed, let $X\in\mathbb{R}^{\ell\times m}$ be a matrix defined by $X(i,i)=1$ for $i=1,\ldots,r$, and $X(i,j)=0$ for all other indices $(i,j)$. Then $X\in\mathcal{F}(\ell,m)$. By choosing some $\hat{X}\in\mathcal{F}(\ell,m)$, let $\hat{R}=A(\mathcal{M})\Pi-A(\mathcal{L})\hat{X}$. We show that $(X,F,G,u)=(\hat{X},\hat{R}^{+},\hat{R}^{-},\|\hat{R}\|_{1})$ is a feasible solution to $\mathsf{P}(\mathcal{L},\mathcal{M})$. It is clear that this tuple satisfies the third, fourth, and fifth constraints. Regarding the first constraint, we have
$A(\mathcal{M})\Pi-A(\mathcal{L})X$ $=A(\mathcal{M})\Pi-A(\mathcal{L})\hat{X}$
$=\hat{R}$
$=\hat{R}^{+}-\hat{R}^{-}$
$=F-G.$
For the second constraint, equality (1) implies that
$\sum_{i=1}^{d}F(i,j)+G(i,j)$ $=\sum_{i=1}^{d}\hat{R}^{+}(i,j)+\hat{R}^{-}(i,j)$
$\leq\|\hat{R}\|_{1}=u$
holds for each $j=1,\ldots,m$. Thus, $\mathsf{P}(\mathcal{L},\mathcal{M})$ has a feasible solution. Accordingly, it follows from the duality theorem that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{M}))$.
Parts (ii)-(iv)  Let $(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{M})$. Then $X^{*}$ is a feasible solution to $\mathsf{H}(\mathcal{L},\mathcal{M})$. Thus, we get $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))\leq\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}$. Furthermore, since $(X^{*},F^{*},G^{*},u^{*})$ satisfies all constraints of $\mathsf{P}(\mathcal{L},\mathcal{M})$, we obtain
$\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}$ $=\|F^{*}-G^{*}\|_{1}$
$=\max_{j=1,\ldots,m}\|F^{*}(:,j)-G^{*}(:,j)\|_{1}$
$\leq\max_{j=1,\ldots,m}\|F^{*}(:,j)\|_{1}+\|G^{*}(:,j)\|_{1}$
$=\max_{j=1,\ldots,m}\sum_{i=1}^{d}F^{*}(i,j)+G^{*}(i,j)$
$\leq u^{*}=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M})).$
Accordingly,
Let $\tilde{X}^{*}$ be the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{M})$, and let $\tilde{R}=A(\mathcal{M})\Pi-A(\mathcal{L})\tilde{X}^{*}$. Then $\tilde{R}$ satisfies the relation $\|\tilde{R}\|_{1}=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$. As shown in part (i), $(X,F,G,u)=(\tilde{X}^{*},\tilde{R}^{+},\tilde{R}^{-},\|\tilde{R}\|_{1})$ is a feasible solution to $\mathsf{P}(\mathcal{L},\mathcal{M})$. Accordingly,
$\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))=\|\tilde{R}\|_{1}\geq\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M})).$ (11)
From (10) and (11), we obtain the equalities,
$\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{M}))$ $=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{M}))$
$=\|A(\mathcal{M})\Pi-A(\mathcal{L})X^{*}\|_{1}=\|\tilde{R}\|_{1}.$
As shown above, $X^{*}$ and $(\tilde{X}^{*},\tilde{R}^{+},\tilde{R}^{-},\|\tilde{R}\|_{1})$ are feasible solutions to $\mathsf{H}(\mathcal{L},\mathcal{M})$ and $\mathsf{P}(\mathcal{L},\mathcal{M})$, respectively. Hence, this completes the proof of parts (ii)-(iv). ∎
Appendix B Proof of Theorem 2
@@
First, we show that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))\geq\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{N}))$ holds under condition (C1). The matrix $X^{*},\Gamma^{*}]$ belongs to $\mathcal{F}(\ell,n)$, and thus, it is a feasible solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$. In addition, we obtain $\|A(\mathcal{L})-A(\mathcal{L})X^{*}\|_{1}=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by Theorem 1 (iii), and $\|A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}]\|_{1}\leq\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by condition (C1) and Theorem 1 (ii). Thus, the objective function value of $\mathsf{H}(\mathcal{L},\mathcal{N})$ at $X^{*},\Gamma^{*}]$ satisfies
@@
Next, we show the reverse inequality. Let $\hat{X}^{*}=[\hat{\bm{x}}_{1}^{*},\ldots,\hat{\bm{x}}_{n}^{*}]$ denote the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$. We find that
@@
The last inequality comes from that the matrix $\hat{\bm{x}}_{1}^{*},\ldots,\hat{\bm{x}}_{\ell}^{*}]$ belongs to $\mathcal{F}(\ell,\ell)$ and is, therefore, a feasible solution to $\mathsf{H}(\mathcal{L},\mathcal{L})$. The two inequalities we showed above yield
@@
We conclude that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{N}))$ and the matrix $X^{*},\Gamma^{*}]$ is the optimal solution to $\mathsf{H}(\mathcal{L},\mathcal{N})$. ∎
@@
  • Lemma 2: Let $\alpha^{*}$ be the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{L})$. If $\alpha^{*}$ satisfies condition (C1), then $\alpha$ can be constructed from $\alpha^{*}$ such that $\alpha$ is a feasible solution to $\mathsf{P}(\mathcal{N},\mathcal{N})$, and the objective function value of $\mathsf{P}(\mathcal{N},\mathcal{N})$ at $\alpha$ is $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$.
  • Lemma 3: Let $\beta^{*}$ be the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$. If $\beta^{*}$ satisfies condition (C2), then $\beta$ can be constructed from $\beta^{*}$ such that $\beta$ is a feasible solution to $\mathsf{D}(\mathcal{N},\mathcal{N})$, and the objective function value of $\mathsf{D}(\mathcal{N},\mathcal{N})$ at $\beta$ is $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))$.
Theorem 1 (i) tells us that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))$. Accordingly, as we reviewed in Section III-B, the weak duality theorem implies that $\alpha$ and $\beta$ are optimal solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$, respectively. Furthermore, it means that $\mathrm{opt}(\mathsf{P}(\mathcal{N},\mathcal{N}))=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. As a result, we can establish part (ii) of Theorem 2.
@@
Let $\mathcal{L}\subset\mathcal{N}$. Let $\alpha^{*}=(X^{*},F^{*},G^{*},u^{*})$ be the optimal solution to $\mathsf{P}(\mathcal{L},\mathcal{L})$. Assume that condition (C1) holds. By letting $\Gamma^{*}=[\bm{\gamma}_{j}^{*}\mid j\in\mathcal{N}\setminus\mathcal{L}]$ for the optimal solution $\bm{\gamma}_{j}^{*}$ to $\mathsf{R}_{j}(\mathcal{L},X^{*})$ and letting $\Pi$ be a permutation matrix of size $n$ satisfying $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$, construct $\alpha=(X,F,G,u)\in\mathbb{R}^{n\times n}\times\mathbb{R}^{d\times n}\times\mathbb{R}^{d\times n}\times\mathbb{R}$ as follows:
@@
where R=AAXR=A-AX for $X$ as constructed above. Then, the following hold.
@@
Part (i)  It is sufficient to prove that $\alpha$ satisfies the first, second, and fifth constraints of $\mathsf{P}(\mathcal{N},\mathcal{N})$, since $F=R^{+}\geq O$ and $G=R^{-}\geq O$. Let us look at the first constraint. Since $R$, $F$ and $G$ are constructed as R=AAXR=A-AX, $F=R^{+}$ and $G=R^{-}$, we have
@@
for $S=A(\mathcal{L})-A(\mathcal{L})X^{*}$ and $T=A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}$. It follows from parts (ii) and (iii) of Theorem 1 that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\|A(\mathcal{L})-A(\mathcal{L})X^{*}\|_{1}=\|S\|_{1}$. It follows from condition (C1) that $\|T\|_{1}=\|A(\mathcal{N}\setminus\mathcal{L})-A(\mathcal{L})\Gamma^{*}\|_{1}\leq\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. Consequently, we obtain $u=\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))$. ∎
@@
Let $\mathcal{L}\subset\mathcal{N}$. Let $\beta^{*}=(Y^{*},Z^{*},\bm{s}^{*},\bm{t}^{*},v^{*})$ be the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$. Assume that condition (C2) holds. By letting $\Pi$ be a permutation matrix of size $n$ satisfying $A\Pi=[A(\mathcal{L}),A(\mathcal{N}\setminus\mathcal{L})]$, and $\Delta=(Y^{*})^{\top}A(\mathcal{N}\setminus\mathcal{L})$, construct $\beta=(Y,Z,\bm{s},\bm{t},v)\in\mathbb{R}^{d\times n}\times\mathbb{R}^{n\times n}\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$ as follows:
@@
[…]
@@
[…]
@@
Let $\alpha$ and $\beta$ be as constructed in Lemmas 2 and 3. From the lemmas and Theorem 1 (i), we find that $\alpha$ and $\beta$ are feasible solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$ and the objective function value of $\mathsf{P}(\mathcal{N},\mathcal{N})$ at $\alpha$ is equal to that of $\mathsf{D}(\mathcal{N},\mathcal{N})$ at $\beta$. Thus, the weak duality theorem implies that $\alpha$ and $\beta$ are optimal solutions to $\mathsf{P}(\mathcal{N},\mathcal{N})$ and $\mathsf{D}(\mathcal{N},\mathcal{N})$. Furthermore, this means that $\mathrm{opt}(\mathsf{P}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{P}(\mathcal{N},\mathcal{N}))$. Consequently, it follows from parts (ii) and (iii) of Theorem 1 that $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{N},\mathcal{N}))$ holds and $X$ of $\alpha$ is the optimal solution to $\mathsf{H}(\mathcal{N},\mathcal{N})$. ∎
Appendix C Proof of Lemma 1
Proof:
We plug in the optimal solution to $\mathsf{D}(\mathcal{L},\mathcal{L})$ for the first constraint and then express the resulting inequality as $M\leq O$ by letting
The trace of $M$ is given as
by $M\leq O$ and $Z^{*}\geq O$. We also have $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))\geq 0$. Indeed, $\mathrm{opt}(\mathsf{D}(\mathcal{L},\mathcal{L}))=\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))$ by parts (i) and (ii) of Theorem 1, and $\mathrm{opt}(\mathsf{H}(\mathcal{L},\mathcal{L}))\geq 0$, since the objective function value of $\mathsf{H}(\mathcal{L},\mathcal{L})$ takes nonnegative values. We thus get
Appendix D Cluster Construction in Step 2-1 of Algorithm 1
We describe the notation used for constructing a cluster $\mathcal{S}_{\ell}$ in step 2-1 of Algorithm 1. We choose a column $\bm{a}_{i}$ of $A$ and sort the columns $\bm{a}_{1},\ldots,\bm{a}_{n}$ of $A$ by their distance to $\bm{a}_{i}$ in ascending order so that